OpenStudy (anonymous):

last one. Using the ratio of perfect squares method, what is square root of 128 rounded to the nearest hundredth?

geerky42 (geerky42):

Just use calculator, What do you get?

OpenStudy (anonymous):

11.31

OpenStudy (anonymous):

thats it?

geerky42 (geerky42):

Yep.

OpenStudy (anonymous):

i get it nowwwww!!!! Yaaay

OpenStudy (anonymous):

its reaaaally easy

geerky42 (geerky42):

Just so you know, this is not how you use "ratio of perfect squares method." But we are not really familiar with it, so yeah.

OpenStudy (anonymous):

oh. crap. i failed my practice session. i got a forty%

geerky42 (geerky42):

just finish math homework and get over it haha

geerky42 (geerky42):

you got right on last two questions though?

OpenStudy (anonymous):

i missed all three

geerky42 (geerky42):

Really? So 11.31 is not right?

OpenStudy (anonymous):

it was 11.30 suuuuuper close

geerky42 (geerky42):

This is so wrong...

OpenStudy (anonymous):

my thing is?

geerky42 (geerky42):

http://www.wolframalpha.com/input/?i=sqrt%28128%29 Must be because of some round up error.

geerky42 (geerky42):

maybe we really need to know this method, sorry for getting you wrong on questions.

OpenStudy (anonymous):

its alright! thanks for trying to help me

OpenStudy (freckles):

@geerky42 I have never heard of that method either and I can not find this method on google.

OpenStudy (freckles):

Perhaps the op can tell us the method.

OpenStudy (freckles):

using the method is probably how they got 11.30 however the method is performed

geerky42 (geerky42):

I think @triciaal is familiar with it, she just did her work on OP's previous question. We would like to know what this method is.

geerky42 (geerky42):

Currently idling lol...

OpenStudy (freckles):

\[11=\sqrt{121}<\sqrt{128}<\sqrt{144}=12\] I know this

OpenStudy (anonymous):

idk man. i just transferred from half a year of pre-al to Algebra senester 2

OpenStudy (anonymous):

Given the equations 9x+34y=6 and 2x+12y=9, by what factor would you multiply the second equation to eliminate y and solve the system through linear combinations? −43 −34 −32 −72 does anyone get this?

OpenStudy (anonymous):

semester*

OpenStudy (anonymous):

dude. make your own thread

geerky42 (geerky42):

Please don't post your question on other user's question. @Nieke

OpenStudy (freckles):

11+12=23 128-121=7 144-128=16 so what is 16/23=.6957 and 7/23=.30435 so somehow I think we suppose (somehow choose) to say 11+.30435 11.30435 which rounded to the nearest hundredth is 11.30

OpenStudy (freckles):

oh by the way 16/23=.6957 and 7/23=.30435 we approximations

OpenStudy (freckles):

I think the method being talked about in this khan video may be the method

geerky42 (geerky42):

Could be it. It fits well.

OpenStudy (anonymous):

did you figure it out? because im not really understanding this khan video